OpenStudy (anonymous):

find the range of f(x)=3-√(2x+5)

OpenStudy (anonymous):

biggest this can be is 3 because \[\sqrt{2x+5}\geq 0\] so \[-\sqrt{2x+5}\leq 0\] and therefore \[3-\sqrt{2x+5}\leq 3\]

OpenStudy (anonymous):

range is \[(-\infty,3]\]